OpenStudy (anonymous):

how many liters each of 15% acid and 33% acid should be mixed to get 120 liters of 21% acid? by using a system of equation I get x=80 and y is equal to 40 is this correct?

hero (hero):

x + y = 120 .15x + .33y = 120(.21) x = 80, y = 40 correct

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

i have one more for you to look at dealing with another system of equattion

hero (hero):

Okay

OpenStudy (anonymous):

a freight train and an express train leave towns 390 km apart, traveling toward one another. The freight train travels 30 km per hr slower than the express train. They pass one another 3 hr later. What are their speeds? use a system of equations

OpenStudy (anonymous):

should i set it up like 390=x+y/30 390=x-y/3 or should i not do that cause then i would get 71.5 for x and 58.5 for y

OpenStudy (anonymous):

so how should i set this one up as a system of equation?

hero (hero):

Hang on a minute, sorry

OpenStudy (anonymous):

oh okay

hero (hero):

We have to use d = rt for this

hero (hero):

We're given that the speed of the express train travels 30km faster than the freight train, so speed of freight train = x speed of express train = x + 30 The distance the express train travels is only part of the total distance. The distance the freight train travels will represent the remaining distance. so let distance of express = y distance of freight = 390 - y Now you can set up the system Freight Train: y = 3x Express Train: 390 - y = 3(x + 30)

hero (hero):

So the system to solve is y = 3x 390 - y = 3(x + 30)

OpenStudy (anonymous):

and from there i get 390-y=3x+90 correct?

hero (hero):

yes, isolate y in the second equation, then use your graphing calc to figure out the intersection point

OpenStudy (anonymous):

then i get 300-y=3x

hero (hero):

You should get something like y = 3x y = 300 - 3x

hero (hero):

Notice now you have two equations which you can graph on your calculator. Or you can just solve manually.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

its not adding up to me

OpenStudy (anonymous):

i think i am supposed to get 80 for x

OpenStudy (anonymous):

and y should be 50

OpenStudy (anonymous):

but the way yo have me doing it gives me 50 for x and 150 for y

OpenStudy (anonymous):

on never mind its supposed to add up to 150 correct?

hero (hero):

The numbers are correct

OpenStudy (anonymous):

okay got it. so i have 50 for x and 150 for y

hero (hero):

Do you know what the numbers mean?

OpenStudy (anonymous):

x is for the speed of the freight train and y is for the speed expressed by the freight train correct

hero (hero):

Yes, what do the numbers 50 and 150 mean?

OpenStudy (anonymous):

x= 50 speed of frieght and y=150 spped of expressed by the freight

hero (hero):

"expressed by the freight" is not the proper way to describe it. There are two trains: one is a freight train, the other is an express train. 50 km/h is the speed of the freight train 150 is the distance the express train travels However, you need to find the speed of the express train.